OpenStudy (anonymous):

sin^2(x) + cos(x) = 2

OpenStudy (shubhamsrg):

sin^2(x) =1-cos^2(x) ..it now forms a quadratic in cos x ..

OpenStudy (anonymous):

I know it becomes (1-cos^2(x)) + cos(x) = 2.

OpenStudy (anonymous):

But I don't know how to solve for x value. The answer should be in degrees or radians.

OpenStudy (callisto):

Not sure if it is correct sin^2(x) + cos(x) = 2 (1-cos^2(x)) + cosx -2=0 cos^2x - cosx +1 =0 => no real solution...

OpenStudy (shubhamsrg):

well it becomes cos^2(x) - cos x + 1 = 0 .. aha..i see..it gives a complex solution..

OpenStudy (saifoo.khan):

-cos^2x + cos x -1 = 0 For problems like these, suppose cos^x = z -z^2 + z -1 = 0 And solve the quadratic.

OpenStudy (shubhamsrg):

the complex solution is not very tough to get here but i guess its of no use here..hmmn..

OpenStudy (apoorvk):

@Callisto you made a 'silly-mistake' in your equation.

OpenStudy (callisto):

Please point it out. It's the nth times I made silly mistakes today. Sorry

OpenStudy (apoorvk):

(1-cos^2(x)) + cosx -2=0 this simplifies to--> -cos^2(x) + cosx -1 = 0 You were doing perfectly otherwise, don't worry.

OpenStudy (callisto):

Divided both sides by -1, wouldn't it be cos^2x - cosx +1 =0?